# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 1 Introduction
Dating to the 1840s, Difference-in-Differences (DiD) is now the most common research design for estimating causal effects in the social sciences. [^1] Currie2020 find that almost 25% of all NBER empirical working papers in 2018 and 17% of empirical articles in economics’ top 5 journals mention them. The earliest DiD applications we are aware of are from Ignaz Semmelweis from the 1840’s (Semmelweis1983) and Snow1855. For a brief overview of the long history of DiD in economics, see Section 2 of Lechner2011. A basic DiD design requires two time periods, one before and one after some treatment begins, and two groups, one that receives a treatment and one that does not. The DiD estimate equals the change in outcomes for the treated group minus the change in outcomes for the untreated group: the difference of two differences. If the average change in the outcomes would have been the same in the two groups had treatment not occurred, a so-called parallel trends assumption, this comparison estimates the average treatment effect among treated units.
@@
For many years, the common practice in applied research was to estimate complex DiD designs using linear regressions with unit and time fixed effects (two-way fixed effects, henceforth TWFE). Their identifying assumptions and interpretation were informally traced to the fact that, in the $2\times 2$ case, a TWFE estimator gives the same estimate as a DiD estimator calculated directly from sample means, and thus inherits a clear causal interpretation under a specific parallel trends identification assumption. This appeared to justify the use of a single technique for any type of design or specification. Recent research, however, has shown that simple regressions can fail to estimate meaningful causal parameters when DiD designs are complex and treatment effects vary, producing estimates that are not only misleading in their magnitudes but potentially of the wrong sign; for recent overviews, see, e.g., Roth2023a, deChaisemartin2023-survey, and Callaway2023_review. The significance of these findings is substantial; given the prevalence of DiD analysis in modern applied econometrics work, common empirical practices have almost certainly yielded misleading results in several concrete cases (baker2022much). [^2] Braghieri_etal_2022_Facebook also shows how newer DiD methods can lead to first-order different results when compared to standard TWFE regressions.
@@
So, what should applied researchers do instead? This paper proposes a unified framework for discussing and conducting DiD studies that is rooted in the principles of causal inference in the presence of treatment effect heterogeneity. The central conclusion of recent methodological research is that even complex DiD studies can be understood as aggregations of $2\times 2$ comparisons between one set of units for whom treatment changes and another set for whom it does not. This fact links a wide variety of DiD designs used in practice and guides methodological choices about estimating them. Viewing DiD studies through the lens of $2\times 2$ “building blocks” aids in interpretability by clarifying that they yield causal quantities that aggregate the treatment effects identified by each $2\times 2$ component. It also means that identification comes from the simple parallel trends assumptions required for each $2\times 2$ building block. Practically, the building block framework suggests first estimating each $2\times 2$ and then aggregating them. As long as the effective sample size is large, this approach allows for asymptotically valid inference using standard techniques.
@@
This framework is a “forward-engineering” approach to DiD that embraces treatment effect heterogeneity and constructs estimators that recover well-motivated causal parameters under explicitly stated assumptions. By fixing the goals of the study (the target parameters) and deriving analytical techniques, forward engineering provides clear benefits over “reverse-engineering” approaches that begin with a familiar regression specification and derive the assumptions under which it has some causal interpretation. The methods we describe in this paper combine familiar techniques with some newer ones, but expressly avoid the difficulties of interpretation inherent in common regression estimators (Goodman2021; deChaisemartin2020; Sun2021; Borusyak2023). Moreover, this interpretation changes across specifications, making it hard to understand the difference between non-robustness and a shifting target parameter. In contrast, our proposed framework naturally leads to estimation procedures that target the same parameter under different transparent identification assumptions. Thus, two estimates can be distinguished easily by their identifying assumptions. Finally, the principles of the forward-engineering approach provide guidance to good econometric practices even in settings without well-established methodological findings.
@@
This paper is not designed to be a comprehensive literature review; its goal is to provide guidelines for practitioners who want to better understand DiD and its various forms. Because of the tremendous variations in design, data, and specification that practitioners encounter, we opt to focus on three of the most common aspects of modern DiD studies: the use of weights, covariates, and staggered treatment timing. We apply techniques to address these issues to a specific example: the causal effect of recent public health insurance expansions in the US on county-level mortality. In an appendix, we briefly discuss related DiD designs with different treatment variables (ones that turn on and off or take many values), additional comparisons (i.e., triple-difference designs), distributional target parameters, or different data structures (repeated cross-sections or unbalanced panels). Several recent reviews follow the logic laid out here and review additional DiD-related topics and technical details Roth2023a; deChaisemartin2023-survey; Callaway2023_review.
@@
The rest of the paper is structured as follows. Section 2 introduces the Medicaid running example. Section 3 discusses the canonical $2\times 2$ DiD setups with and without weights, and Section 4 discusses threats to the identification assumptions, how to assess them, and how to incorporate covariates to make them more plausible, all of this still in a $2\times 2$ setup. Section 5 extends the $2\times 2$ setup to multiple periods with potentially staggered treatment adoption. Section 6 concludes and briefly discusses some extensions that involve more complex DiD designs.
@@ 2 Medicaid and Mortality: The running example
  • Notes: The table shows which states adopted the ACA’s Medicaid expansion in each year as well as the share of all states, counties, and adults in each expansion year.
@@
States also expanded Medicaid largely because of economic and political considerations (sommers2013us), creating observable differences between expansion and non-expansion states. For instance, just four out of the 22 states that expanded Medicaid in 2014 are in the southern Census region compared to seven out of ten non-expansion states. This suggests a potential role for covariates when analyzing Medicaid expansion.
@@
Finally, mortality is measured in jurisdictions like states and counties, which are of very different sizes. Choices about (population) weights not only determine how different estimation approaches average the units within a given expansion group but also how a given estimation technique averages estimated effects across those groups. California, for example, represented 4.5 percent of the states that expanded Medicaid in 2014, 5 percent of the counties, but 23.1 percent of the adults ages 20-64; its contribution to “the” average outcome for the 2014 expansion group is very different with weights than without. The final three columns of Table 1 show that in our data the entire 2014 expansion group contains 44 percent of the states, 36 percent of the counties, but 45 percent of all adults. Weighting will, therefore, change how important the estimated treatment effects are for the 2014 group.
@@
Several recent papers study the effect of ACA Medicaid expansion on mortality rates for lower-income adults who are most likely to gain insurance through Medicaid. miller2021medicaid and wyse2024 use simple DiD methods to provide credible evidence that Medicaid reduced adult mortality rates for targeted sub-populations. Unfortunately, their analyses require restricted links between income and mortality data, which are important to overcome the low statistical power in studies using aggregate mortality data (black2022simulated). Our goal is to pursue a replicable and shareable example based on a related analysis by Borgschulte2020. They use a sophisticated strategy to select and use covariates in a weighted TWFE regression using restricted access data, and find that Medicaid expansion reduced aggregate county-level mortality rates. We use only publicly available data, which allows us to share a fully-reproducible replication package, and consider only a handful of intuitive demographic and economic covariates sufficient to illustrate several practical challenges that can arise with DiD. This empirical exercise is meant solely to illustrate how to tackle several common features of DiD designs. The results are pedagogical in spirit and do not represent the best possible estimates of Medicaid’s effect on adult mortality.
@@
Our outcome variable is the crude adult mortality rate, $Y_{i,t}$, for people ages 20-64 (measured per 100,000) by county ($i$) from 2009 to 2019 released by the CDC_Vital_Statistics. [^3] It is common to adjust mortality rates by the county age distribution. Unfortunately, the CDC measurements of age-specific deaths for many counties are restricted due to there being fewer than ten annual deaths. We aim to use publicly available and shareable data for pedagogical purposes; we follow Borgschulte2020 and use the crude mortality rate. We denote county $i$’s adult population in 2013 as $W_{i}$ and its socioeconomic covariates in year $t$ (discussed below) as $X_{i,t}$. The information in Table 1 defines the treatment group variable $G_{i}$ that equals the year in which county $i$’s state expanded Medicaid with $G_{i}=\infty$ for the non-expansion states. Our final sample contains 2,604 counties in states with complete data on mortality rates from 2009 to 2019 and covariates for 2013 and 2014.
@@
Faced with a setup such as this, researchers need to make a range of tightly related choices. Which treatment groups in Table 1 should be compared to each other and over what time horizons? What must be true for those comparisons to identify causal effects, and how should one empirically evaluate their plausibility? How can other information, such as covariates or pre-period outcomes, be used to improve the credibility of the design? How do these methodological choices affect the causal interpretation of a given analysis? The aim of this review is to demonstrate to practitioners using DiD in realistic scenarios why and how to use state-of-the-art econometric tools to answer these questions.
@@
We begin our discussion by focusing on the canonical $2\times 2$ DiD setup, which has two time periods—one before and one after treatment—and two groups—one that remains untreated in both periods and one that becomes treated in the second period. In our Medicaid example, we focus on comparisons between the 2014 expansion group (978 counties) and the non-expansion group as of 2019 (1,222 counties) in 2014 and 2013. When we consider more complex designs, this kind of comparison will still play a role: it will be one $2\times 2$ “building block” among many.
@@ 3.1 Causal effects and target parameters: the ATT
The first step of any causal analysis is to define the causal quantity of interest, also called the target parameter. We use the potential outcomes framework of Rubin1974 and Robins1986 to do so. Let $Y_{i,t}(0,0)$ denote unit $i$’s potential outcome at time $t$ if it remained untreated in both periods. Analogously, let $Y_{i,t}(0,1)$ denote unit $i$’s potential outcome at time $t$ if untreated in the first period but exposed to treatment by the second period. In our example, $Y_{i,t}(0,0)$ is county $i$’s mortality rate in period $t$ in a world in which Medicaid did not expand in its state, and $Y_{i,t}(0,1)$ is its mortality rate in a world in which Medicaid did expand in 2014.444The stable unit treatment value assumption requires that the only treatment determining county $i$’s mortality rate is its own. Medicaid expansion in neighboring counties must not affect deaths in county $i$. Without this assumption, potential outcomes need to account for many different counterfactuals as they are a function of many treatment variables. To simplify notation, we will write $Y_{i,t}(0)=Y_{i,t}(0,0)$ and $Y_{i,t}(1)=Y_{i,t}(0,1)$, as the potential outcomes are defined by treatment exposure in period two (Medicaid expansion status by 2014). Nonetheless, it will be useful for later discussions that these potential outcomes correspond to treatment sequences.
@@
We additionally assume that county mortality rates were not affected by the Medicaid expansion before Medicaid expanded, which will be crucial for justifying the structure of the DiD estimator (see, e.g., Abbring2003, Malani2015, and Roth2023a). This standard “no anticipation” assumption ensures that we observe untreated potential outcomes before Medicaid expansion takes effect: $Y_{i,2013}=Y_{i,2013}(0)$. It also helps us define effective treatment dates. For instance, if the announcement of Medicaid expansion affects mortality before its actual expansion, “treatment” begins when the policy is announced rather than implemented. We formally state this assumption for completeness and maintain it throughout the paper.
@@
The potential outcomes define a causal effect for every unit in every time period, $Y_{i,t}(1)-Y_{i,t}(0)$. These describe what Medicaid expansion did to mortality rates in a specific treated county or what it would have done in a specific untreated county. This framework allows for arbitrary heterogeneity in the effects across units and time, i.e., the effect of Medicaid expansion can be different in every county and year. But it is hard to learn about this degree of rich heterogeneity without additional strong restrictions. Instead, DiD analyses typically seek to estimate averages of heterogeneous treatment effects. In particular, most DiD designs target the average treatment effect on the treated at time $t$, or $ATT(t)$:
@@
$ATT(t)$ compares (weighted) average observed post-expansion mortality rates among treated counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}|D_{i}=1]$) to the (weighted) average untreated mortality rates for the same treated counties (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}(0)|D_{i}=1]}$). The second quantity is counterfactual because untreated outcomes are never observed for treated counties. Note that, by the no-anticipation assumption, $ATT(t)=0$ for all pre-treatment periods, i.e., $ATT(2013)=0$ in our two-period Medicaid example. This ensures that cross-group outcome comparisons before treatment begins reflect untreated potential outcome gaps, which is central to the logic of DiD. Note that we abuse notation and omit the weight index when defining $ATT$’s; we do that to unclutter notation throughout the paper.
Equation (3.1) shows that weighting enters the analysis early on, as part of the definition of the causal parameter. If interest lies in the average treatment effect of Medicaid on mortality in the average treated county, sometimes motivated by a view of jurisdictions as “laboratories of democracy” none to be prioritized over another, then the relevant target parameter is an equally weighted average ($\omega=1$). If, on the other hand, the parameter of interest is the average treatment effect of Medicaid on mortality in the county in which the average treated adult lives, then population weights are appropriate ($\omega=W$). When treatment effect heterogeneity is related to the weights, weighted and unweighted target parameters differ meaningfully (solon2015). We conduct some of our empirical exercises with and without population weights to highlight how weights affect a given DiD result. From a policy-relevancy perspective, we argue that using population weights in our Medicaid expansion application is probably more appropriate.
In the Medicaid context, the unweighted $ATT(2014)$ answers the question, “What was the average causal effect of Medicaid expansion on 2014 mortality rates among the 2014 expansion state counties?” Whether this parameter (or any causal parameter) is “of interest” is an argument about theoretical importance, policy relevance, and how one is planning to use it. Other target parameters are also possible. Designs other than DiD identify different kinds of average treatment effects, and some DiD methods use quantile (Athey2006; Callaway2019) or distribution regression (Fernandez-Val2024) approaches to target features of the marginal distributions of $Y_{i,t}(1)$ and $Y_{i,t}(0)$ among treated units. We focus on identification and estimation strategies that target ATT parameters but emphasize that the $2\times 2$ building block framework applies to DiD methods more broadly; see our appendix for more discussions about distributional parameters.

### v2

@@ 1 Introduction
Dating to the 1840s, difference-in-differences (DiD) is now the most common research design for estimating causal effects in the social sciences. [^1] Currie2020 find that almost 25% of all NBER empirical working papers and 17% of empirical articles in five leading general-interest economics journals in 2018 mention DiD. The earliest DiD applications we are aware of are from Ignaz Semmelweis from the 1840s (Semmelweis1983) and Snow1855. For a brief overview of the long history of DiD in economics, see Section 2 of Lechner2011. A basic DiD design requires two time periods, one before and one after some treatment begins, and two groups, one that receives a treatment and one that does not. The DiD estimate equals the change in outcomes for the treated group minus the change in outcomes for the untreated group: the difference of two differences. If the average change in the outcomes would have been the same in the two groups had the treatment not occurred, which is referred to as a “parallel trends” assumption, this comparison estimates the average treatment effect among treated units.
@@
For many years, the common practice in applied research was to estimate complex DiD designs using linear regressions with unit and time fixed effects (two-way fixed effects, henceforth TWFE). Their identifying assumptions and interpretation were informally traced to the fact that, in the $2\times 2$ case, a TWFE estimator gives the same estimate as a DiD estimator calculated directly from sample means, and thus inherits a clear causal interpretation under a specific parallel trends identification assumption. This appeared to justify the use of a single technique for any type of design or specification. Recent research, however, has shown that simple regressions can fail to estimate meaningful causal parameters when DiD designs are complex and treatment effects vary, producing estimates that are not only misleading in their magnitudes but potentially of the wrong sign. The significance of these findings is substantial; given the prevalence of DiD analysis in modern applied econometrics work, common empirical practices have almost certainly yielded misleading results in several concrete cases (baker2022much).
@@
So, what should applied researchers do instead? This paper proposes a unified framework for discussing and conducting DiD studies that is rooted in the principles of causal inference in the presence of treatment effect heterogeneity. The central conclusion of recent methodological research is that even complex DiD studies can be understood as aggregations of $2\times 2$ comparisons between one set of units for which treatment changes and another set for which it does not. This fact links a wide variety of DiD designs used in practice and guides methodological choices about estimating them. Viewing DiD studies through the lens of $2\times 2$ “building blocks” aids in interpretability by clarifying that they yield causal quantities that aggregate the treatment effects identified by each $2\times 2$ component. It also means that identification comes from the simple parallel trends assumptions required for each $2\times 2$ building block. Practically, this framework suggests first estimating each $2\times 2$ building block and then aggregating them. As long as the effective sample size is large, this approach allows for asymptotically valid inference using standard techniques.
@@
This framework is a “forward-engineering” approach to DiD that embraces treatment effect heterogeneity and constructs estimators that recover well-motivated causal parameters under explicitly stated assumptions. By fixing the goals of the study (the target parameters) and deriving analytical techniques, forward engineering provides clear benefits over “reverse-engineering” approaches that begin with a familiar regression specification and derive the assumptions under which it has some causal interpretation. The methods we describe in this paper combine familiar techniques with some newer ones, but expressly avoid the difficulties of interpretation inherent in common regression estimators (Goodman2021; deChaisemartin2020; Sun2021; Borusyak2023). Moreover, the interpretation of common regression estimators changes across specifications, which makes it hard to understand the difference between non-robustness and a shifting target parameter. In contrast, our proposed framework naturally leads to estimation procedures that target the same parameter under different transparent identification assumptions. Thus, two estimates can be distinguished easily by their identifying assumptions. Finally, the principles of the forward-engineering approach provide guidance to good econometric practices even in settings without well-established methodological findings.
@@
This paper is not designed to be a comprehensive literature review; its goal is to provide guidelines for practitioners who want to better understand DiD and its various forms. Because of the tremendous variations in design, data, and specification that practitioners encounter, we opt to focus on three of the most common aspects of modern DiD studies: the use of weights, covariates, and staggered treatment timing. We apply techniques to address these issues to a specific example: the causal effect of recent public health insurance expansions in the US on county-level mortality. In an appendix, we briefly discuss related DiD designs with different treatment variables (ones that turn on and off or take many values), additional comparisons (i.e., triple-difference designs), distributional target parameters, or different data structures (repeated cross-sections or unbalanced panels). Several recent reviews follow the logic laid out here and cover additional DiD-related topics and technical details:Roth2023a; deChaisemartin2023-survey; Callaway2023_review.
@@
The rest of the paper is structured as follows. Section 2 introduces the Medicaid example. Section 3 discusses the canonical $2\times 2$ DiD setups with and without weights, and Section 4 discusses threats to the identification assumptions, how to assess them, and how to incorporate covariates. Section 5 extends the $2\times 2$ setup to multiple periods with potentially staggered treatment adoption. Section 6 concludes and briefly discusses some extensions that involve more complex DiD designs.
@@ 2 Medicaid and mortality: The running example
  • The table shows which states adopted the ACA’s Medicaid expansion in each year as well as the share of all states, counties, and adults in each expansion year.
@@
States expanded Medicaid largely because of economic and political considerations (sommers2013us), which created observable differences between expansion and non-expansion states. For instance, just four out of the 22 states that expanded Medicaid in 2014 are in the southern Census region; conversely, seven out of 10 non-expansion states are in the South. This suggests a potential role for covariates when analyzing Medicaid expansion.
@@
Finally, mortality is measured in jurisdictions like states and counties, which are of very different sizes. Choices about (population) weights determine not only how different estimation approaches average the units within a given expansion group but also how a given estimation technique averages estimated effects across those groups. California, for example, represented 4.5% of the states that expanded Medicaid in 2014, 5.4% of the counties, but 27.7% of the adults ages 20-64; its contribution to “the” average outcome for the 2014 expansion group is very different with weights than without. The final three columns of Table 1 show that, in our data the entire 2014 expansion group contains 44% of the states, 36% of the counties, but 45% of all adults. Weighting will therefore change how important the estimated treatment effects are for the 2014 group.
@@
Several recent papers study the effect of ACA Medicaid expansion on mortality rates for lower-income adults, who are most likely to gain insurance through Medicaid. miller2021medicaid and wyse2024 use simple DiD methods to provide evidence that Medicaid reduced adult mortality rates for targeted sub-populations. Unfortunately, their analyses require restricted links between income and mortality data, which are important for overcoming the low statistical power in studies using aggregate mortality data (black2022simulated). Our goal is to pursue a replicable and shareable example based on a related analysis by Borgschulte2020. They use a sophisticated strategy to select and use covariates in a weighted TWFE regression using restricted access data, and find that Medicaid expansion reduced aggregate county-level mortality rates. We use publicly available data, which we include in a fully-reproducible replication package, and consider only a handful of intuitive demographic and economic covariates sufficient to illustrate several practical challenges that can arise with DiD. This empirical exercise is meant solely to illustrate how to tackle several common features of DiD designs. The results are pedagogical in spirit and do not represent the best possible estimates of Medicaid’s effect on adult mortality.
@@
Our outcome variable is the crude adult mortality rate, $Y_{i,t}$, for people ages 20-64 (measured per 100,000) by county ($i$) from 2009 to 2019 released by the CDC_Vital_Statistics. [^2] It is common to adjust mortality rates by the county age distribution. Unfortunately, the CDC measurements of age-specific deaths are restricted for counties with fewer than 10 annual deaths. We aim to use publicly available and shareable data for pedagogical purposes; we follow Borgschulte2020 and use the crude mortality rate. We denote county $i$’s adult population in 2013 by $W_{i}$ and its socioeconomic covariates in year $t$ (discussed below) by $X_{i,t}$. The information in Table 1 defines the treatment group variable $G_{i}$ that equals the year in which county $i$’s state expanded Medicaid; $G_{i}=\infty$ for the non-expansion states. Our final sample contains 2,604 counties in states with complete data on mortality rates from 2009 to 2019 and covariates for 2013 and 2014.
@@
Faced with a setup such as this, researchers need to make a range of tightly related choices. Which treatment groups in Table 1 should be compared with each other and over what time horizons? What must be true for those comparisons to identify causal effects, and how should one empirically evaluate their plausibility? How can other information, such as covariates or pre-period outcomes, be used to improve the credibility of the design? How do these methodological choices affect the causal interpretation of a given analysis? The aim of this review is to demonstrate to practitioners using DiD in realistic scenarios why and how to use state-of-the-art econometric tools to answer these questions.
@@
We begin our discussion by focusing on the canonical $2\times 2$ DiD setup, which has two time periods—one before and one after treatment—and two groups—one that remains untreated in both periods and one that becomes treated in the second period. In our Medicaid example, we focus on comparisons in 2014 and 2013 between the 2014 expansion group (978 counties) and the group that had not expanded by 2019 (1,222 counties). When we consider more complex designs, this kind of comparison will still play a role: it will be one $2\times 2$ “building block” among many.
@@ 3.1 Causal effects and target parameters: The ATT
The first step of any causal analysis is to define the causal quantity of interest, also called the target parameter. We use the potential outcomes framework of Rubin1974 and Robins1986 to do so. Let $Y_{i,t}(0,0)$ denote unit $i$’s potential outcome at time $t$ if it remained untreated in both periods. Analogously, let $Y_{i,t}(0,1)$ denote unit $i$’s potential outcome at time $t$ if untreated in the first period but exposed to treatment by the second period. In our example, $Y_{i,t}(0,0)$ is county $i$’s mortality rate in period $t$ in a world in which Medicaid did not expand in its state, and $Y_{i,t}(0,1)$ is its mortality rate in a world in which Medicaid did expand in 2014.333We have implicitly introduced the stable unit treatment value assumption, which holds if the only treatment determining county $i$’s mortality rate is its own. In other words, Medicaid expansion in neighboring counties must not affect deaths in county $i$. If this fails, then there are effectively many different treatment variables and counterfactuals. The potential outcomes notation and ensuing analysis would then need to account for this. To simplify notation, we will write $Y_{i,t}(0)=Y_{i,t}(0,0)$ and $Y_{i,t}(1)=Y_{i,t}(0,1)$, as the potential outcomes are defined by treatment exposure in period two (Medicaid expansion status by 2014). Nonetheless, it will be useful for later discussions that these potential outcomes correspond to treatment sequences.
@@
We additionally assume that county mortality rates were not affected by the Medicaid expansion before Medicaid expanded, which is crucial to the validity of the DiD estimator (see, e.g., Abbring2003, Malani2015, and Roth2023a). This standard “no anticipation” assumption ensures that we observe untreated potential outcomes before Medicaid expansion takes effect: $Y_{i,2013}=Y_{i,2013}(0)$. It also helps us define effective treatment dates. For instance, if the announcement of Medicaid expansion affected mortality before its actual expansion, “treatment” would begin when the policy was announced rather than implemented. We formally state this assumption for completeness and maintain it throughout the paper.
@@
The potential outcomes define a causal effect for every unit in every time period, $Y_{i,t}(1)-Y_{i,t}(0)$. These describe what Medicaid expansion did to mortality rates in a specific treated county or what it would have done in a specific untreated county. This framework allows for arbitrary heterogeneity in the effects across units and time; that is, it allows the effect of Medicaid expansion to be different in every county and year. But it is hard to learn about this degree of rich heterogeneity without additional strong conditions holding. Instead, DiD analyses typically seek to estimate (weighted) averages of heterogeneous treatment effects. In particular, most DiD designs target the average treatment effect on the treated at time $t$, or $ATT(t)$:
@@
Equation (3.1) shows that $ATT(t)$ compares (weighted) average observed post-expansion mortality rates among treated counties ($\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}|D_{i}=1]$) with the (weighted) average untreated mortality rates for the same treated counties (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathop{}\!\mathbb{E}_{\omega}[Y_{i,t}(0)|D_{i}=1]}$). The second quantity is counterfactual because untreated outcomes are never observed for treated counties. Note that by the no-anticipation assumption, $ATT(t)=0$ for all pre-treatment periods; that is, $ATT(2013)=0$ in our two-period Medicaid example. This ensures that cross-group outcome comparisons before treatment begins reflect untreated potential outcome gaps, which is central to the logic of DiD. Note that we abuse notation and omit the weight index when defining $ATT$’s; we do that to unclutter notation throughout the paper.
Equation (3.1) shows that weighting enters the analysis early on, as part of the definition of the causal parameter. In the Medicaid context, the unweighted $ATT(2014)$ answers the question, “What was the average causal effect of Medicaid expansion on 2014 mortality rates among the 2014 expansion state counties?” The weighted $ATT(2014)$ answers the question, “What was the average causal effect of Medicaid expansion on 2014 mortality rates among adults in counties in states that expanded Medicaid in 2014?” This point interacts with other justifications for weighting, such as improving precision. With heterogeneous treatment effects, adopting a weighting scheme designed to improve precision in the presence of heteroskedasticity in a constant-coefficient regression model will also change the target parameter, potentially by a lot when treatment effects are correlated with the weights (solon2015). Comparing weighted and unweighted estimates, therefore, does not show whether weighting matters for estimation or inference; these reflect different target parameters. In our example, the population-weighted parameter is probably more policy-relevant, but we conduct some of our empirical exercises both ways to show how weighting can affect a given DiD result. Which parameter is “of interest” is an argument about theoretical importance, policy relevance, and the use to which it will be put.
Other target parameters are also possible. Designs other than DiD identify different kinds of average treatment effects, and some DiD methods use quantile regression (Athey2006; Callaway2019) or distribution regression (Fernandez-Val2024) approaches to target features of the marginal distributions of $Y_{i,t}(1)$ and $Y_{i,t}(0)$ among treated units. We focus on identification and estimation strategies that target ATT parameters but emphasize that the $2\times 2$ building block framework applies to DiD methods more broadly; see our appendix for more discussions about distributional parameters.